### v1

@@ 4. Proof of main theorems
$2\|\nabla_{\!\text{A}}\psi_{R}\|^{2}+2\Im\int x_{k}B_{jk}\psi_{R}^{*}\partial_{j}^{A}\psi_{R}+\Re\langle 2x\cdot\nabla_{\!\text{A}}\psi_{R}+d\psi_{R},-\tfrac{i}{2}(\boldsymbol{\alpha}\cdot B\cdot\boldsymbol{\alpha})\psi_{R}\rangle\\ -\langle\psi_{R},(x\cdot\nabla)(V^{2})\psi_{R}\rangle-m\langle\psi_{R},(x\cdot\nabla\{\beta,V\})\psi_{R}\rangle\\ +\Re\langle 2x\cdot\nabla_{\!\text{A}}\psi_{R}+d\psi_{R},-i(\{\boldsymbol{\alpha},V\}\cdot\nabla_{\!\text{A}})\psi_{R}\rangle+\Re\langle 2x\cdot\nabla_{\!\text{A}}\psi_{R}+d\psi_{R},-i(\boldsymbol{\alpha}\cdot\nabla V)\psi_{R}\rangle\\ =\Re\langle 2x\cdot\nabla_{\!\text{A}}\psi_{R}+d\psi_{R},(H_{m}(A,V)+\lambda)-i(\boldsymbol{\alpha}\cdot\nabla\xi_{R})\psi\rangle.$ (4.2)
@@
$\left(2-\left(\frac{4d-6}{d-2}\right)\varepsilon_{1}-\left(\frac{4}{d-2}\right)\varepsilon_{2}\,\varepsilon_{3}-2m\varepsilon_{4}^{2}-\left(\frac{8d-8}{d-2}\right)\varepsilon_{3}-\left(\frac{4d-4}{d-2}\right)\varepsilon_{2}\right)\|\nabla_{\!\text{A}}\psi_{R}\|^{2}\\ \leq\Re\langle 2(x\cdot\nabla_{\!\text{A}})\psi_{R}+d\psi_{R},(H_{m}(A,V)+\lambda)-i(\boldsymbol{\alpha}\cdot\nabla)\xi_{R}\psi\rangle$ (4.3)
@@ Acknowledgements
There is no data associated with this article
@@
There is no conflict of interest to declare

### v2

@@ 4. Proof of main theorems
$2\|\nabla_{\!\text{A}}\psi_{R}\|^{2}+2\Im\int x_{k}B_{jk}\psi_{R}^{*}\partial_{j}^{A}\psi_{R}+\Re\langle 2x\cdot\nabla_{\!\text{A}}\psi_{R}+d\psi_{R},-\tfrac{i}{2}(\boldsymbol{\alpha}\cdot B\cdot\boldsymbol{\alpha})\psi_{R}\rangle\\ -\langle\psi_{R},(x\cdot\nabla)(V^{2})\psi_{R}\rangle-m\langle\psi_{R},(x\cdot\nabla\{\beta,V\})\psi_{R}\rangle\\ +\Re\langle 2x\cdot\nabla_{\!\text{A}}\psi_{R}+d\psi_{R},-i(\{\boldsymbol{\alpha},V\}\cdot\nabla_{\!\text{A}})\psi_{R}\rangle+\Re\langle 2x\cdot\nabla_{\!\text{A}}\psi_{R}+d\psi_{R},-i(\boldsymbol{\alpha}\cdot\nabla V)\psi_{R}\rangle\\ =\Re\langle 2x\cdot\nabla_{\!\text{A}}\psi_{R}+d\psi_{R},(H_{m}(A,V)+\lambda-i(\boldsymbol{\alpha}\cdot\nabla\xi_{R}))\psi\rangle.$ (4.2)
@@
$\left(2-\left(\frac{4d-6}{d-2}\right)\varepsilon_{1}-\left(\frac{4}{d-2}\right)\varepsilon_{2}\,\varepsilon_{3}-2m\varepsilon_{4}^{2}-\left(\frac{8d-8}{d-2}\right)\varepsilon_{3}-\left(\frac{4d-4}{d-2}\right)\varepsilon_{2}\right)\|\nabla_{\!\text{A}}\psi_{R}\|^{2}\\ \leq\Re\langle 2(x\cdot\nabla_{\!\text{A}})\psi_{R}+d\psi_{R},(H_{m}(A,V)+\lambda-i(\boldsymbol{\alpha}\cdot\nabla)\xi_{R})\psi\rangle$ (4.3)
@@
Since the arguments needed to prove Theorem 1.8 are similar to the ones in Theorem 1.2, we will present its proof next, for clarity’s sake.
Proof of Theorem 1.8.
As in the proof of Theorem 1.2, we assume by contradiction that $\psi\in H^{1}(\mathbb{R}^{d})$ is a nontrivial solution of the eigenvalue equation (3.3), for which we can derive in the same way as before the equation (4.2). Moreover, in order to have an estimate involving the eigenvalue $\lambda$, we need, in this case, an additional identity: multiplying equation (3.5) by $\psi_{R}$ and integrating, then taking the real part of the resulting identity we get
$\|\nabla_{A}\psi_{R}\|^{2}+\Re\langle\psi_{R},-\tfrac{i}{2}(\boldsymbol{\alpha}\cdot B\cdot\boldsymbol{\alpha})\psi_{R}\rangle+m^{2}\|\psi_{R}\|^{2}+\|V\psi_{R}\|^{2}\\ +\Re\langle\psi_{R},m\{\beta,V\}\psi_{R}\rangle+\Re\langle\psi_{R},i(\{\boldsymbol{\alpha},V\}\cdot\nabla_{\!\text{A}})\psi_{R}\rangle-\Re\langle\psi_{R},i(\boldsymbol{\alpha}\cdot\nabla V)\psi_{R}\rangle\\ =\lambda^{2}\|\psi_{R}\|^{2}+\Re\langle\psi_{R},(H_{m}(A,V)+\lambda-i(\boldsymbol{\alpha}\cdot\nabla\xi_{R}))\psi\rangle.$ (4.4)
Now, taking the difference between (4.2) and (4.4), that is taking (4.2) - (4.4) gives
$\|\nabla_{A}\psi_{R}\|^{2}+2\Im\int x_{k}B_{jk}\psi_{R}^{*}\partial_{j}^{A}\psi_{R}+\Re\langle 2x\cdot\nabla_{\!\text{A}}\psi_{R}+d\psi_{R},-\tfrac{i}{2}(\boldsymbol{\alpha}\cdot B\cdot\boldsymbol{\alpha})\psi_{R}\rangle\\ -\langle\psi_{R},(x\cdot\nabla)(V^{2})\psi_{R}\rangle-m\langle\psi_{R},(x\cdot\nabla\{\beta,V\})\psi_{R}\rangle\\ +\Re\langle 2x\cdot\nabla_{\!\text{A}}\psi_{R}+d\psi_{R},-i(\{\boldsymbol{\alpha},V\}\cdot\nabla_{\!\text{A}})\psi_{R}\rangle+\Re\langle 2x\cdot\nabla_{\!\text{A}}\psi_{R}+d\psi_{R},-i(\boldsymbol{\alpha}\cdot\nabla V)\psi_{R}\rangle\\ -\Re\langle\psi_{R},-\tfrac{i}{2}(\boldsymbol{\alpha}\cdot B\cdot\boldsymbol{\alpha})\psi_{R}\rangle-m^{2}\|\psi_{R}\|^{2}-\|V\psi_{R}\|^{2}\\ -\Re\langle\psi_{R},m\{\beta,V\}\psi_{R}\rangle-\Re\langle\psi_{R},i(\{\boldsymbol{\alpha},V\}\cdot\nabla_{\!\text{A}})\psi_{R}\rangle+\Re\langle\psi_{R},i(\boldsymbol{\alpha}\cdot\nabla V)\psi_{R}\rangle\\ =-\lambda^{2}\|\psi_{R}\|^{2}-\Re\langle\psi_{R},(H_{m}(A,V)+\lambda-i(\boldsymbol{\alpha}\cdot\nabla\xi_{R}))\psi\rangle\\ +\Re\langle 2x\cdot\nabla_{\!\text{A}}\psi_{R}+d\psi_{R},(H_{m}(A,V)+\lambda-i(\boldsymbol{\alpha}\cdot\nabla\xi_{R}))\psi\rangle.$ (4.5)
Now we shall estimate each term of (4.5). Using assumption (1.24) one has
$2\left|\Im\int x_{k}B_{jk}\psi_{R}^{*}\partial_{j}^{A}\psi_{R}\right|$ $\leq 2\int|x||B||\psi_{R}||\nabla_{\!\text{A}}\psi_{R}|$
$\leq\frac{1}{\varepsilon_{0}}\int|x|^{2}|B|^{2}|\psi_{R}|^{2}+\varepsilon_{0}\|\nabla_{A}\psi_{R}\|^{2}$
$\leq\left(\frac{\varepsilon_{1}}{\varepsilon_{0}}+\varepsilon_{0}\right)\|\nabla_{A}\psi_{R}\|^{2}+\frac{\delta}{\varepsilon_{0}}\|\psi_{R}\|^{2}.$
For the second term in the first line of (4.5) we have
[…]
The first summand in the right-hand side is bounded in the same way as before. For the second, by using the Hardy inequality, we get
$\frac{d}{2}\int|B||\psi_{R}|^{2}$ $=\frac{d}{2}\int|x||B||\psi_{R}|\frac{|\psi_{R}|}{|x|}$
$\leq\frac{d}{(d-2)}\left(\int|x|^{2}|B|^{2}|\psi_{R}|^{2}\right)^{1/2}\|\nabla_{\!\text{A}}\psi_{R}\|$
$\leq\frac{d}{2(d-2)}\left[\left(\frac{\varepsilon_{1}}{\varepsilon_{0}}+\varepsilon_{0}\right)\|\nabla_{A}\psi_{R}\|^{2}+\frac{\delta}{\varepsilon_{0}}\|\psi_{R}\|^{2}\right].$
Similarly, one can estimate the first term in the fourth line of (4.5). The terms in the second and third line of (4.5) can be treated as in Theorem 1.2. Therefore it is left to estimates the remaining terms that depend on $V$ and that come from the introduction of the new identity (4.4).
Using (1.7) and Hardy’s inequality, we get
$\|V\psi_{R}\|^{2}\leq\varepsilon_{3}^{2}\int\frac{|\psi_{R}|^{2}}{|x|^{2}}\leq\frac{4\varepsilon_{3}^{2}}{(d-2)^{2}}\|\nabla_{A}\psi_{R}\|^{2}.$
Similarly
$|\Re\langle\psi_{R},m\{\beta,V\}\psi_{R}\rangle|$ $\leq 2m\int|V||\psi_{R}|^{2}$
$\leq 4m\frac{\varepsilon_{3}}{d-2}\|\nabla_{A}\psi_{R}\|\|\psi_{R}\|$
$\leq 4m^{2}\frac{\varepsilon_{3}^{2}}{(d-2)^{2}}\|\nabla_{A}\psi_{R}\|^{2}+\|\psi_{R}\|^{2}.$
Again using (1.7), one has
$|\Re\langle\psi_{R},i(\{\boldsymbol{\alpha},V\}\cdot\nabla_{\!\text{A}})\psi_{R}\rangle|$ $\leq 2\int|V||\psi_{R}||\nabla_{A}\psi_{R}|$
$\leq\frac{4\varepsilon_{3}}{(d-2)}\|\nabla_{A}\psi_{R}\|^{2}.$
Finally
$|\Re\langle\psi_{R},i(\boldsymbol{\alpha}\cdot\nabla V)\psi_{R}\rangle|$ $\leq\int|\nabla V||\psi_{R}|^{2}$
$\leq\frac{2\varepsilon_{2}}{(d-2)}\|\nabla_{A}\psi_{R}\|^{2}.$
Using these estimates in (4.5) and taking the limit as $R$ goes to infinity (using Lemma 3.2) we obtain
$\left(1-\frac{(4d-5)}{(d-2)}\left(\frac{\varepsilon_{1}}{\varepsilon_{0}}+\varepsilon_{0}\right)-\left(\frac{4}{d-2}\right)\varepsilon_{2}\,\varepsilon_{3}-2m\varepsilon_{4}^{2}-\left(\frac{8d-6}{d-2}\right)\varepsilon_{3}\right.\\ \left.-\left(\frac{4d-2}{d-2}\right)\varepsilon_{2}-\frac{4}{(d-2)^{2}}\varepsilon_{3}^{2}-\frac{4m^{2}}{(d-2)^{2}}\varepsilon_{3}^{2}\right)\|\nabla_{A}\psi\|^{2}+\left(\lambda^{2}-m^{2}-1-\frac{\delta}{\varepsilon_{0}}\frac{(4d-5)}{2(d-2)}\right)\|\psi\|^{2}\leq 0.$ (4.6)
In virtue of (1.23) and (1.25) the left hand side of (4.6) is non negative, and consequently $\psi=0$, which gives us a contradiction. ∎
@@ Acknowledgements
There is no data associated with this article.
@@
There is no conflict of interest to declare.